1. A bullet of density is 5.40 (g)/(cc)and a mass of 80.0 g is dropped into a 100.0 ml graduated cylinder containing exactly 50.0 ml of water. To what height will the water rise in the cylinder?

2. Three hundred fifty seven grams of zinc is submerged into a graduated cylinder filled with 130 cc of water. The water level rises to 180 cc reading. What is the density of zinc?
1
Expert's answer
2017-11-09T04:47:07-0500
1. As far as we know, a submerged object displaces a volume of fluid equal to its own volume. Then, simply summing up the volume of the water in cylinder and the volume of a bullet, we will get the level of the water after the immersion of a bullet. Thus, let’s start with calculation of the volume of the bullet:
V=m/d= (80 (g))/(5.40 (g cm^(-3)))=14.8 cm^3=14.8 ml
The new height of water in the cylinder is:
V^'= 50.0 ml+14.8 ml=64.8 ml
2. This is an inverse problem. While we know the starting and the final levels of water, we can get the volume of the zinc piece and its density. The volume of zinc submerged is:
V_Zn=180cm^3-130 cm^3=50 cm^3
The density of zinc is:
d_Zn=m/V=357(g)/50(cm^3 ) =7.14 g/cm^3

Answer: 1. 64.8 ml; 2. 7.14 g/cm^3.